OpenStudy (anonymous):

need help (steps) for 4-(3x+1)-6x=5(x+4)

OpenStudy (ajprincess):

4-(3x+1)-6x=5(x+4) 4-3x-1-6x=5x+20 Can u do nw?

OpenStudy (anonymous):

yes ,thanks

OpenStudy (ajprincess):

yw.:)